OpenStudy (anonymous):

(secx)dy/dx=e^(y+sinx), please help me solve the differential equation. Thanks :)

OpenStudy (anonymous):

move secx to the other side

OpenStudy (anonymous):

then moved dx to the other side, then integrate it. THAT"S ALL

zepdrix (zepdrix):

\[\large \sec x\frac{dy}{dx}=e^{(y+\sin x)}\] Recall a rule of exponents:\[\large e^{a+b}=e^a\cdot e^b\]

zepdrix (zepdrix):

\[\large \sec x\frac{dy}{dx}=e^{(y+\sin x)} \qquad \rightarrow \qquad \large \sec x\frac{dy}{dx}=e^{y}\cdot e^{\sin x}\]So that would be our first step right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

and then it would be \[e^y/secx \times e^{sinx}/secx\]

OpenStudy (abb0t):

First order Seperable differential equation. You can think of it as "cross multiplying" Even though that's what it seems like, its not. but you could think of it as doing just that. You're getting dy w/ y's and dx with x's.

zepdrix (zepdrix):

Yes. Next move the e^y to the side containing dy.

zepdrix (zepdrix):

Woops the way you wrote it joker, it looks like your denominator is sec^2x. Careful. There should only be 1 secx on the bottom on the right side of our equation.

OpenStudy (anonymous):

so \[\frac{ 1 }{ e^y }dy=e^{sinx}/secxdx\]?

zepdrix (zepdrix):

\[\huge e^{-y}dy\qquad =\qquad \frac{e^{\sin x}}{\sec x}dx\] Yes very good. We can simplify the 1/sec x. Remember what that will give us in terms of sines or cosines?

OpenStudy (anonymous):

\[e^{-y}dy=e ^{sinx}times cosx\]

OpenStudy (anonymous):

1/secx becomes cosx

zepdrix (zepdrix):

\[\huge e^{-y}dy\qquad =\qquad (\cos x) e^{\sin x}dx\] Ok good :) From here you can integrate.

OpenStudy (anonymous):

do I use natural log?

zepdrix (zepdrix):

You will use a `U substitution` to get through these integrals. If your problem is asking for an `explicit` solution, then yes we'll have to use natural log after we integrate.

OpenStudy (abb0t):

\[\frac{ 1 }{ \sec(x) } = \cos(x)\]

OpenStudy (anonymous):

so then its u=sinx, du=cosx dx

OpenStudy (anonymous):

\[e^{-y} dy=e^udu\]?

zepdrix (zepdrix):

k looks good so far.

OpenStudy (anonymous):

after taking the log \[-ydy=udu\]

zepdrix (zepdrix):

No don't take the log right now. Integrate first.

OpenStudy (anonymous):

after integrating \[-e^{-y}=e^u\]

zepdrix (zepdrix):

\[\large -e^{-y}=e^u+C\] Ok good.

zepdrix (zepdrix):

Make sure to move that negative over before applying the natural log.

OpenStudy (anonymous):

so then it becomes \[e^{-y}=-e^u ?\]

zepdrix (zepdrix):

Don't forget your C, it's quite important in this problem.

zepdrix (zepdrix):

\[\large e^{-y}=C-e^u\]

OpenStudy (anonymous):

\[-y=C-u\]

OpenStudy (anonymous):

so is the final answer \[y=sinx+c\] ?

zepdrix (zepdrix):

Taking the natural log of both sides,\[\large \ln(e^{-y})=\ln(C-e^u)\]Giving us,\[\large -y=\ln(C-e^u) \qquad \rightarrow \qquad y=-\ln(C-e^{u})\]

zepdrix (zepdrix):

Im not sure what happened to your natural log on the right side. It doesn't disappear like it does on the left.

OpenStudy (anonymous):

oh, right, forgot the natural log on the other side. :) Thanks so much zepdrix! :)

zepdrix (zepdrix):

np c: